The Cellular Connection charges $60 for a phone and $29 per month under its economy plan. Write an equation that can be used to determine the total cost,
C(t), of operating a Cellular Connection phone for t months. Then find the total cost for 6 months.

Answer by richwmiller(17219)

y=29x+60
let x=6
y=29*6+60 for a six month plan
y=6*(29+10)=6*39